OpenStudy (anonymous):

What is the equation of the line that passes through the points (0, 5) and (2, 3)?

OpenStudy (igreen):

First find the slope: Slope Equation: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) In this case: \(y_2 = 3\) \(y_1 = 5\) \(x_2 = 2\) \(x_1 = 0\) So: \(m = \dfrac{3-5}{2-0}\) \(m = \dfrac{-2}{2}\) \(m = -1\)\ So your slope is -1. We can write it in the form of y = mx + b Where \(m\) is slope and where \(b\) is the y-intercept. We already have the y-intercept. (0, 5) And we have the slope which is -1. So: y = -1x + 5 or y = -x + 5

OpenStudy (igreen):

@Dequavia123

OpenStudy (anonymous):

yesss

OpenStudy (igreen):

@Dequavia123 I solved it for you.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'.